Write the equation of a line that passes through (4,1) and is parallel to 5x+y=2
Answer by mananth(16949)

5 x + 1 y = 2
Find the slope of this line
make y the subject
1 y = -5 x + 2
Divide by 1
y = -5 x + 2
Compare this equation with y=mx+b
slope m = -5
The slope of a line parallel to the above line will be the same
The slope of the required line will be -5
m=-5,point(4,1)
Find b by plugging the values of m & the point in
y=mx+b
1=-20 +b
b= 21
m= -5
Plug value of the slope and b in y = mx +b
The required equation is y=-5 x +21

m.ananth@hotmail.ca